Analyzing Manufacturing Cost Accounts

Fire Rock Company manufactures designer paddle boards in a wide variety of sizes and styles. The following incomplete ledger accounts refer to transactions that are summarized for June:

Materials
June 1 Balance 33,000 June 30 Requisitions (A)
June 30 Purchases 132,700


Work in Process
June 1 Balance (B) June 30 Completed jobs (F)
June 30 Materials (C)
June 30 Direct labor (D)
June 30 Factory overhead applied (E)


Finished Goods
June 1 Balance 0 June 30 Cost of goods sold (G)
June 30 Completed jobs (F)


Wages Payable
June 30 Wages incurred 125,000


Factory Overhead
June 1 Balance 21,000 June 30 Factory overhead applied (E)
June 30 Indirect labor (H)
June 30 Indirect materials 17,700
June 30 Other overhead 89,800

In addition, the following information is available:

a. Materials and direct labor were applied to six jobs in July:

Job No. Style Quantity Direct Materials Direct Labor
201 T100 190 $20,020 $14,000
202 T200 390 36,950 25,000
203 T400 210 14,880 9,000
204 S200 310 41,210 32,000
205 T300 180 22,660 17,000
206 S100 100 5,500 3,000
Total 1,380 $141,220 $100,000

b. Factory overhead is applied to each job at a rate of 150% of direct labor cost.

c. The June 1 Work in Process balance consisted of two jobs, as follows:

Job No. Style Work in Process, June 1
201 T100 $5,400
202 T200 15,600
Total $21,000

d. Customer jobs completed and units sold in July were as follows:

Job No. Style Completed in July Units Sold in July
201 T100 X 152
202 T200 X 312
203 T400 0
204 S200 X 260
205 T300 X 149
206 S100 0

1. Determine the missing amounts associated with each letter and complete the following table. If required, round amounts to the nearest dollar. If an answer is zero, enter in "0". Enter all amounts as positive numbers.

Job No. Quantity June 1
Work in
Process
Direct
Materials
Direct
Labor
Factory
Overhead
Total Cost Unit Cost Units Sold Cost of Goods Sold
No. 201 $ 5,400 $ 20,020 $ 14,000 $ $ $ $
No. 202 15,600 36,950 25,000
No. 203 14,880 9,000
No. 204 41,210 32,000
No. 205 22,660 17,000
No. 206 5,500 3,000
Total $21,000 141,220 100,000 $ $ $

a. Materials Requisitions $

b. Work in Process Beginning Balance $

c. Direct Materials $

d. Direct Labor $

e. Factory overhead applied $

f. Completed jobs $

g. Cost of goods sold $

h. Indirect labor $

2. Determine the June 30 balances for each of the inventory accounts and factory overhead. Use the minus sign to indicate any credit balances.

Materials: $
Work in Process: $
Finished Goods: $
Factory Overhead: $
Answer #1

1.

Job No. Quantity June 1 Work in Process Direct Material Direct labor Factory Overhead Total Cost per unit Units Sold Cost of Goods Sold
201 190 $            5,400 $           20,020 $           14,000 $           21,000 $          60,420 $          318.00 152 $          48,336
202 390 $          15,600 $           36,950 $           25,000 $           37,500 $       115,050 $          295.00 312 $          92,040
203 210 $           14,880 $             9,000 $           13,500 $          37,380 $          26,700
204 310 $           41,210 $           32,000 $           48,000 $       121,210 $          391.00 260 $        101,660
205 180 $           22,660 $           17,000 $           25,500 $          65,160 $          362.00 149 $          53,938
206 100 $             5,500 $             3,000 $             4,500 $          13,000
Totals 1380 $          21,000 $         141,220 $         100,000 $         150,000 $       412,220 $        322,674

a. Materials Requisitions = $141220+17700 = $158920

b. Work in Process Beginning Balance = $21000

c. Direct Materials = $141220

d. Direct Labor = $100000

e. Factory overhead applied = $150000

f. Completed jobs = $60420+115050+121210+65160 = $361840

g. Cost of goods sold = $322674

h. Indirect labor = $125000-100000 = $25000

2.
Material = $33000+132700-158920 = $6780
Work in Process = $37380+13000 = $50380
Finished Goods = $361840-322674 = $39166
Factory Overhead = $17700+25000+21000+89800-150000 = $3500